# Gaia Barium Dwarfs and Their Ostensibly Ordinary Counterparts

### v1

@@ Appendix B Additional Plots
Figure 8 shows Ba abundance as a function of [Fe/H] and M as a complementary plot to figure 3 of R24. Y abundance plots that complement Figure 2 are presented in Figures 9 10.
@@ Appendix C Data Tables
Gaia DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] |
3036788173176962048 | – | ||||||
4790658375100436736 | |||||||
3208829887146528384 | |||||||
6091100873778091392 | |||||||
5620705644518714752 | |||||||
5406806102703951360 | |||||||
3772123820203876736 | – | ||||||
5397347794427206656 | |||||||
6181668466886465280 | |||||||
2914005812780342272 | – | ||||||
5722712938857848192 | – | ||||||
5164360455867611264 | – | ||||||
2923820637686349568 | – | ||||||
4795585179328518016 | |||||||
5550946678313327744 | – | ||||||
6169594184243050752 | – | ||||||
5327240562221030528 | – | ||||||
2964151136511557120 | – | ||||||
5317828124219671168 | – | ||||||
6300835156308377344 | |||||||
5329433194583181952 | – | ||||||
2919286664051008640 | – | ||||||
6192821814055372800 | – | ||||||
5744310130364616576 | – | ||||||
3697174751104247808 | – | ||||||
5890631694717361408 | – | ||||||
3147042801161881472 | – | ||||||
2923409729576118912 | – | ||||||
5091887624392418816 | – | ||||||
3633619813626022784 |
Gaia DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] |
6472197475176313472 | |||||||
6645994082722928768 | |||||||
4878036323641744384 | |||||||
4646881493008180352 | |||||||
4774266821393569792 | |||||||
6489348928855798272 | |||||||
5495572631335262208 | |||||||
5707071630039205248 | |||||||
4715941990147018752 | |||||||
4881348087024892416 | |||||||
5363208336349003648 | |||||||
2436150255390344320 | |||||||
3623765616756619392 | |||||||
6669412928776138112 | |||||||
4693571910684108928 | |||||||
6081919436288801024 | |||||||
5435770125183875200 | |||||||
2600643791975931264 | |||||||
2545472433093014912 | |||||||
4650039354829199616 | |||||||
3285029929897944320 | |||||||
4622751301587564416 | |||||||
5371201888966580992 | |||||||
5386048560107201792 | |||||||
6402712399045698560 |

### v2

@@ Appendix B Additional Plots
Figure 9 shows Ba abundance as a function of [Fe/H] and M as a complementary plot to figure 3 of R24. Y abundance plots that complement Figure 2 are presented in Figures 9 and 10.
@@ Appendix C Data Tables
| Gaia DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] | T | |||
| 3036788173176962048 | – | ||||||||||
| 4790658375100436736 | |||||||||||
| 3208829887146528384 | |||||||||||
| 5620705644518714752 | |||||||||||
| 6091100873778091392 | |||||||||||
| 5406806102703951360 | |||||||||||
| 3772123820203876736 | – | ||||||||||
| 5397347794427206656 | |||||||||||
| 6181668466886465280 | |||||||||||
| 5722712938857848192 | – | ||||||||||
| 2914005812780342272 | – | ||||||||||
| 5164360455867611264 | – | ||||||||||
| 4795585179328518016 | |||||||||||
| 5550946678313327744 | – | ||||||||||
| 2923820637686349568 | – | ||||||||||
| 5327240562221030528 | – | ||||||||||
| 5329433194583181952 | – | ||||||||||
| 2964151136511557120 | – | ||||||||||
| 6169594184243050752 | – | ||||||||||
| 6300835156308377344 | |||||||||||
| 5317828124219671168 | – | ||||||||||
| 2919286664051008640 | – | ||||||||||
| 5744310130364616576 | – | ||||||||||
| 5890631694717361408 | – | ||||||||||
| 3697174751104247808 | – | ||||||||||
| 6192821814055372800 | – | ||||||||||
| 3147042801161881472 | – | ||||||||||
| 2923409729576118912 | – | ||||||||||
| 5091887624392418816 | – | ||||||||||
| 3633619813626022784 |
Note. — This table is also available in machine-readable format in the arXiv ancillary material with uncertainties provided in separate columns. The quoted error for M1 represents the mean of the upper and lower uncertainties. Quantities above are rounded to 2 decimal digits for readability.
| Gaia DR3 ID | GALAH DR3 ID | Period (d) | Eccentricity | M () | M1 () | [Fe/H] | [Ba/Fe] | [Y/Fe] |
| 6472197475176313472 | 170514003301152 | |||||||
| 6645994082722928768 | 170712004201185 | |||||||
| 4878036323641744384 | 150112002501291 | |||||||
| 4646881493008180352 | 140809004901046 | |||||||
| 4774266821393569792 | 161218003101346 | |||||||
| 6489348928855798272 | 170911004201207 | |||||||
| 5495572631335262208 | 171205004601232 | |||||||
| 5707071630039205248 | 170417002601038 | |||||||
| 4715941990147018752 | 171001002901085 | |||||||
| 4881348087024892416 | 171106003601092 | |||||||
| 5363208336349003648 | 140311007101125 | |||||||
| 2436150255390344320 | 161105003601033 | |||||||
| 3623765616756619392 | 160602001601256 | |||||||
| 6669412928776138112 | 160521004801173 | |||||||
| 4693571910684108928 | 140809004901240 | |||||||
| 6081919436288801024 | 170508003301353 | |||||||
| 5435770125183875200 | 151225003801120 | |||||||
| 2600643791975931264 | 140710006102184 | |||||||
| 2545472433093014912 | 131217001801154 | |||||||
| 4650039354829199616 | 161219003601130 | |||||||
| 3285029929897944320 | 141104004301120 | |||||||
| 4622751301587564416 | 160919004601297 | |||||||
| 5371201888966580992 | 160524002101389 | |||||||
| 5386048560107201792 | 150408004101125 | |||||||
| 6402712399045698560 | 170614005101079 |
Note. — The full table from R24 is available in machine-readable format in the arXiv ancillary material with uncertainties provided in separate columns. The quoted error for M1 represents the mean of the upper and lower uncertainties. Quantities above are rounded to 2 decimal digits for readability.